### v1

@@ 3.3.3 Spatial interaction and Cross-Modulation
The Spatial Interaction and Cross-Modulation level within MASAG dynamically enhances segmentation by weaving together the spatially refined inputs and (derived from Equation 3). Here, with its refined local detail, is enriched by global contexts from , through spatial weights applied via a sigmoid, creating . Conversely, , initially embodying weighted broader contexts, incorporates detailed contexts from , evolving into . This mutual enhancement ensures that both and integrate a comprehensive mix of local and global contexts, solidified by fusing and via multiplication, integrates detailed precision with overarching insights, effectively segmenting complex structural variations. The formulation of spatial interaction is described as follows:
@@ 4.4 Qualitative Results
We showcase visual segmentation results for the Synapse (Figure 2) and ISIC (Figure 3) datasets to further demonstrate the capacity of our method. It is evident that the predictions closely match the actual anatomical structure of the organs in the Synapse dataset maintaining detailed features at the edges, precisely identifying organ areas. Furthermore, a visual assessment of the ISIC dataset shows that the boundaries predicted by our proposed model closely align with the ground truth, surpassing earlier methods, particularly in complex cases characterized by intricate color distribution and irregular borders, owing to our boundary-aware model and the MASAG module, which ensures optimal feature fusion.
@@ 4.5 Ablation Study
We conduct an ablation study on the Synapse dataset to examine the effectiveness of our modules under various configurations as summarized in Table 3. Specifically, the inclusion of MASAG (last three rows) demonstrates its capability to fine-tune focus adaptively, emphasizing key areas and dynamically adjusting features to enhance segmentation, evidenced by improved DSC and reduced HD scores. Besides, the final row demonstrates that our design elements work together effectively, enhancing segmentation predictions by leveraging the synergistic effects of all components for superior detail preservation and precise organ delineation.
@@ 5 Conclusion
We proposed a novel medical image segmentation network to alleviate the restriction of cohort studies in local, global information processing. Our key idea is to mitigate the limitations of fixed receptive sizes and basic fusion techniques in skip connections, thereby enhancing medical image segmentation challenged by the variability in sizes and shapes of anatomical structures. We achieved this through a novel MASAG including multi-scale fusion, spatial selection, interaction, cross-modulation, and recalibration. Our MSANet shows robust performance across two benchmark datasets compared to SOTA methods. Ablation studies confirm the effectiveness of the design and underlying assumptions of our approach.

### v2

@@ 3.3.3 Spatial interaction and Cross-Modulation
The Spatial Interaction and Cross-Modulation level within MASAG dynamically enhances segmentation by weaving together the spatially refined inputs and (derived from Equation 3). Here, with its refined local detail, is enriched by global contexts from , through spatial weights applied via a sigmoid, creating . Conversely, , initially embodying weighted broader contexts, incorporates detailed contexts from , evolving into . This mutual enhancement ensures that both and integrate a comprehensive mix of local and global contexts, solidified by fusing and via multiplication, this integration combines detailed precision with overarching insights, effectively segmenting complex structural variations. The formulation of spatial interaction is described as follows:
@@ 4.4 Qualitative Results
We showcase visual segmentation results for the Synapse (Figure 2) and ISIC (Figure 3) datasets to further demonstrate the capacity of our method. It is evident that the predictions closely match the actual anatomical structure of the organs in the Synapse dataset, maintaining detailed features at the edges and precisely identifying organ areas. Furthermore, a visual assessment of the ISIC dataset shows that the boundaries predicted by our proposed model closely align with the ground truth, surpassing earlier methods, particularly in complex cases characterized by intricate color distribution and irregular borders, owing to our boundary-aware model and the MASAG module, which ensures optimal feature fusion.
@@ 4.5 Ablation Study
We conducted an ablation study on the Synapse dataset to examine the effectiveness of our modules under various configurations as summarized in Table 3. Specifically, the inclusion of MASAG (last three rows) demonstrates its capability to fine-tune focus adaptively, emphasizing key areas and dynamically adjusting features to enhance segmentation, as evidenced by improved DSC and reduced HD scores. Besides, the final row demonstrates that our design elements work together effectively, enhancing segmentation predictions by leveraging the synergistic effects of all components for superior detail preservation and precise organ delineation.
@@ 5 Conclusion
We proposed a novel medical image segmentation network to alleviate the restrictions of cohort studies in local and global information processing. Our key idea is to mitigate the limitations of fixed receptive sizes and basic fusion techniques in skip connections, thereby enhancing medical image segmentation, which is challenged by the variability in sizes and shapes of anatomical structures. We achieved this through a novel MASAG, which includes multi-scale fusion, spatial selection, interaction, cross-modulation, and recalibration. Our MSANet shows robust performance across two benchmark datasets compared to SOTA methods. Ablation studies confirm the effectiveness of the design and underlying assumptions of our approach.